C-8 Landscaping- At an excavation site a conical pile of earth 9 ft high and 14 ft in diameter at the base must be hauled away by a small truck with a capacity of 2.6 cu yd. How many truckloads will it take to transport the soil?
Answer by checkley71(8403)

V=1/3*PI*R^2*H
V=(3.14*7^2*9)/3
V=1,384.74/3
V=461.58 CUBIC FT OF SAND.
461.58/2.6*27 [THE 27 IS THE NUMBER OF CUBIC FEET IN A CUBIC YARD.]
461.58/70.2
6.57 TRUCK LOADS NEEDED TO REMOVE THE SAND.